OpenStudy (anonymous):

sqrt-6 * sqrt-18 Answers: 6sqrt3 -6sqrt3 6i sqrt3 -6i sqrt3

OpenStudy (hba):

\[what \ is \ \sqrt{-1}=\]

OpenStudy (hba):

@girlinthemirror Answer ?

OpenStudy (anonymous):

\[\sqrt{6}\iota * \sqrt{18}\iota\] \[\sqrt{6}\iota * \sqrt{6}i \sqrt{3}i\] \[6i^{2}\sqrt{3}i\] \[-6\sqrt{3}\iota\]

OpenStudy (anonymous):

ie. None of the above o.o Unless I'm wrong in keeping the i's in... Yeh thats wrong XD The answer should be -6sqrt(3)

OpenStudy (hba):

@ScottB05 Giving direct answers is not allowed. Please check the CoC- http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

That's not a direct answer; that's working in how to do the question with the answer??? is it not O.o

OpenStudy (anonymous):

that's working and*

OpenStudy (anonymous):

Thank you @ScottB05 for your help

OpenStudy (anonymous):

So when you come to tackle a similar type of question you can make reference to that and apply the same method therefore learning from it. If I quote the CoC "and not just give them an answer.". With respect @hba I did not give just an answer; but infact a method of approach that can be followed. however I'll take in onboard and next time be less discreet =)

OpenStudy (hba):

@ScottB05 As a rule of thumb, we would prefer the askers to participate in the solution process instead of having the answerers doing all the work. thats not really our concern; we are aimed at being a study site, not a free answering service. We would rather promote learning the material as opposed to just handing over answers. This social contract is aimed at curbing people from using the site to cheat with ...

OpenStudy (anonymous):

That's no problem. The CoC in that case is rather vague =( I misunderstood the rules and I'm sorry - I was wrong, but it won't happen in the future - Thanks for clearing it up thought =)

OpenStudy (hba):

Thanks a lot for understanding :)